OpenStudy (anonymous):

A triangle drawn on a map has sides of lengths 7.0 cm, 11 cm, and 13 cm. The shortest of the corresponding real-life distances is 91 km. Find the longest of the real-life distances. Round your answer to the nearest whole number.

OpenStudy (anonymous):

7:91 = 13:x 7/91 = 13/x 7x = 1183 x = 169 So, the longest of real life distances is 169 km.